Genetics help!
hat the mutant QUESTION 2 to search
Answer #1

Answer

1)

A) all of the offsprings have barred eyes.

Explanation : if the mutant alle is dominant then they will suppress the white allele which is recessive to it and hence all offspring will be barred.

2) D)

Explanation : cross between

AaBbCCddEeff x AaBbCcDDEEff

Consider with monohybrid cross for each individual alleles.

Aa x Aa

- AA=1/4

-Aa=1/2

-aa= 1/4

Bb x Bb

- BB=1/4

-Bb=1/2

-bb=1/4

CC x Cc

- CC=1/2

- Cc=1/2

dd x DD

- Dd = all

Ee x EE

-EE=1/2

-Ee=1/2

ff x ff

- ff= all

So probability of AABBCCDdEeff will be

1/4 * 1/4 * 1/2 * 1 * 1/2 * 1= 1/64

3) A) 3/4